(3.2.21) A sample of blood pressure measurements is taken for a group of​ adults, and those values​ (mm Hg) are listed below. The values are matched so that 10 subjects each have a systolic and diastolic measurement. Find the coefficient of variation for each of the two​ samples; then compare the variation.
Systolic 119,128,157,94,157,121,118,136,125,122
Diastolic 79,76,75,54,89,88,60,65,73,83,
1
Expert's answer
2020-11-16T17:44:40-0500

"CV=\frac{standard deviation}{mean}×100"

Systolic mean "=\frac{1277}{10}=127.7"

Systolic standard deviation is 18.7856

Diastolic mean"=\frac{742}{10}=74.2"

Diastolic standard deviation is 11.5931

Systolic CV "=\frac{18.7856}{127.7}×100=14.71%"

Diastolic CV"=\frac{11.5941}{74.2}×100=15.62%"

The coefficients of variation are relatively equal, however, Diastolic CV is higher by less than 1%. The difference is small, hence, the two variables have approximately the same variation from the mean.